OpenStudy (anonymous):

Use elimination or substitutuion to solve 4x+4y=46 and x+3y=14.5

OpenStudy (mathstudent55):

Let's use elimination: 4x + 4y = 46 x + 3y = 14.5 Multiply the bottom equation by -4 and write below the first equation. Then add them together. 4x + 4y = 46 -4x - 12y = -58 ------------------- 0 - 8y = -12 -8y = -12 y = -12/-8 y = 1.5 Now substitute y = 1.5 in the second original equation: x + 3(1.5) = 14.5 x + 4.5 = 14.5 x = 10 Solution: x = 10, y = 1.5